OpenStudy (anonymous):

Mr. Daton purchased men't sports coats for $80 each and sold them for $158 each. Find the markup on retail to the nearest whole percent.

OpenStudy (campbell_st):

the percentage is \[percentage=\frac{(158 - 80)}{80} \times 100 \]

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

@campbell_st The answer isn't coming out right for some reason.. It is supposed to be 49%.

OpenStudy (campbell_st):

oops I gave you mark up on cost. For markup on retail markup = retail price - cost price = 158 - 80 = $78 percentage = 78/158 * 100 = 49.3%